OpenStudy (anonymous):

Evaluate the following expression. 6^-3

OpenStudy (anonymous):

Is that an exponent or? Inputting what you put into wolframalpha.com I got 1/216

geerky42 (geerky42):

\(\Large 6^{-3} = \dfrac{1}{6^3} = \boxed{\dfrac{1}{216}}\)